# UniADC: A Unified Framework for Anomaly Detection and Classification

### v1

@@ 3.4 Multi-task Discriminator
We employ a pre-trained vision encoder to extract multi-scale image features $\{x_{1},x_{2},\cdots,x_{N}\}=\varphi(X)$, and fuse these features into a fine-grained representation through a feature fusion network $f=\theta(x_{1},x_{2},\cdots,x_{N})$, where $f\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C}$. The fusion network follows a fully convolutional architecture that progressively fuses intermediate-layer features from high to low levels. We then employ a text encoder to embed the names of anomaly categories to obtain $\{g_{1},g_{2},\cdots,g_{Y}\}$, where each $g_{y}\in\mathbb{R}^{C}$ denotes the embedding of the $y$-th anomaly category. The similarity between the anomaly category embedding $g_{y}$ and the feature map $f$ at a spatial location $(h,w)$ is computed as:
@@
where $\tau$ represents the anomaly score threshold. We use the anomaly category with the largest pixel area in $S_{c}$ as the image-level anomaly classification result $I_{c}$. If all pixels in $S_{c}$ are classified as normal, then $I_{c}$ is set to 0. We optimize the feature fusion network using Binary Focal loss [21] and Dice loss [24] for anomaly detection, and Cross-Entropy loss for anomaly classification. The overall loss function is:
@@
where $\cdot$$\cdot$] denotes the concatenation operation, and $\lambda$ is a weight hyperparameter. The normal regions in $M_{c}$ are ignored during the computation of $\mathcal{L}_{CE}$. Unlike prior works [14, 20], we discard normal prompts to avoid performance degradation caused by the severe class imbalance between normal pixels and per-category anomalous pixels.
4 Experiments
4.1 Experimental Setup
Figure 4: Qualitative results of UniADC on the MVTec-FS dataset under $\text{zero-shot}_{\>(K_{n}=2)}$ and $\text{few-shot}_{\>(K_{n}=2,\,K_{a}=1)}$ settings.
@@ 4.1 Experimental Setup
Datasets. We conduct extensive evaluations on three anomaly detection and classification datasets, including MVTec-FS [23], MTD [12], and WFDD [3]. The MVTec-FS dataset is introduced by ZipAF [23] as an extension of the MVTec-AD benchmark [1] for anomaly classification. It contains 15 types of industrial products, each with an average of four anomaly categories. The MTD dataset [12] consists of 1,344 images of magnetic tiles, including five types of anomalies. The subtle differences among these anomalies make the classification task particularly challenging. The WFDD dataset [3] contains 4,101 images across four types of fabrics. We reclassified the dataset by grouping defects of the same type, resulting in an average of three anomaly types per fabric. More details of the WFDD dataset [3] are provided in the Appendix. Some popular anomaly detection datasets, such as VisA [50], are excluded from this work due to the lack of anomaly category labels.
@@
Implementation Details. We adopt a consistent experimental setup for both anomaly prior-guided and anomaly sample-guided inpainting. Specifically, we use the DDIM sampler [31] with 1,000 original diffusion steps. The noise factor $\gamma$ is uniformly sampled between 0.4 and 0.6, and the number of accelerated sampling steps is set to 10. For mask generation, we leverage BiRefNet [47] for binary segmentation to ensure that the mask is located within the foreground region. We set the mini-batch size to 32 for category consistency selection, and generate 16 samples with a resolution of $512\times 512$ per anomaly category for discriminator training. For the multi-task discriminator, we adopt two experimental settings, namely UniADC(CLIP) and UniADC(DINO), which use the CLIP ViT-L/14 [25] and the DINOv3-based dino.txt [30] as the vision-language backbone of the discriminator, respectively. We set both the anomaly score threshold $\tau$ and the loss weight $\lambda$ to 0.5. The list of anomaly prior is provided in the Appendix.
@@
Metrics. We use the Area Under the Receiver Operating Characteristic Curve (I-AUC) metric to evaluate image-level anomaly detection performance. For pixel-level anomaly location, we use Pixel-AUROC (P-AUC) and Per Region Overlap (PRO) [2] metrics. In addition, we report the Top-1 classification accuracy (Acc) and Mean Intersection over Union (mIoU) to evaluate image-level and pixel-level anomaly classification performance, respectively. For the “combined” category in MVTec-FS dataset [23], we only evaluate its anomaly detection performance and exclude it from the anomaly classification metrics, as in previous works [23, 13].
@@
Baselines. In the zero-shot setting, we adopt InCTRL [49], PatchCore [27], PromptAD [20], AnomalyGPT [8], and WinCLIP [14] as baseline methods. For AnomalyGPT [8], we cast classification as a single-choice QA task by prompting the model with candidate anomaly types. For WinCLIP [14], we provide it with anomaly category descriptions and compute the similarity between patch features and anomaly category embeddings, obtaining the anomaly classification results. In the few-shot setting, we combine anomaly detection methods such as PatchCore [27], PromptAD [20], WinCLIP [14], and AnomalyCLIP [48] with the anomaly classification method ZipAF [23] as our baseline methods. We employ the threshold selection strategy proposed in AnomalyNCD [13] to determine the anomaly score threshold, and fine-tune PromptAD [20] and AnomalyCLIP [48] using few-shot anomaly samples to enhance their anomaly detection performance.
4.2 Experimental Results
Figure 5: Examples of synthetic anomaly samples generated by UniADC on the MVTec-FS dataset.
@@ 4.2 Experimental Results
Tables 1 and 2 report the anomaly detection and classification results under zero-shot and few-shot settings, respectively. In the zero-shot setting, most baselines struggle to identify specific anomaly categories due to the absence of real anomaly samples. Although some methods incorporate Vision-Language Models (VLMs) or Large Language Models (LLMs) with anomaly priors, they still fail to achieve satisfactory anomaly detection and classification performance. In contrast, UniADC fully exploits the relevance between anomaly detection and classification, thereby achieving promising performance in both tasks. Compared to other methods, it achieves approximately a 20% improvement in classification accuracy and a 10% improvement in mIoU. Moreover, UniADC(DINO) achieves superior detection and classification accuracy over UniADC(CLIP), highlighting the remarkable performance of DINOv3 [30] in extracting fine-grained image features. In the few-shot setting, UniADC delivers substantial performance gains, even with only one anomaly sample per category, demonstrating its low cost and effective data utilization. Furthermore, it markedly outperforms other two-stage methods, showcasing the benefit of unifying anomaly detection and classification. Qualitative results in Fig. 4 further illustrate its ability to accurately localize and classify anomalous regions under various settings, emphasizing its strong generalization capability and practical applicability.
In addition, UniADC can be extended to the setting of full-shot normal samples, with the results provided in Table 3. We use the unsupervised anomaly detection methods PatchCore [27], RD4AD [4], and RealNet [46], as well as the semi-supervised method BGAD [41], in combination with the anomaly classification method ZipAF [23] as our baselines. When sufficient normal samples are available, UniADC can achieve anomaly detection performance comparable to mainstream full-shot methods. Moreover, it significantly improves anomaly classification performance, addressing the limitations of existing approaches in this aspect and demonstrating its unique practical value.
@@
Table 3: Comparison of UniADC with alternative methods on the MVTec-FS dataset using full-shot normal samples, with $K_{a}=1$.
@@
Fig. 5 presents synthetic anomaly samples generated by UniADC. By incorporating an inpainting control network and reliable category consistency selection strategies, UniADC can generate high-quality anomaly samples that are mask-consistent, category-aligned, and visually diverse. These properties enable UniADC to generalize well across a wide range of anomaly detection and classification scenarios. A comparison with alternative anomaly synthesis methods is provided in the Appendix.
4.3 Ablation Study
We conduct extensive experiments to evaluate the effectiveness of the UniADC components. By default, all ablation experiments use DINOv3 [30] as the backbone of the discriminator. Table 4 investigates the impact of anomaly classification loss and category consistency selection on the performance of UniADC. When the anomaly classification loss is removed, UniADC degenerates into a standard anomaly detection model. In the zero-shot setting, image-level anomaly detection performance improves slightly, while anomaly localization performance degrades. In the few-shot setting, all metrics decline, indicating that incorporating the anomaly classification task does not compromise anomaly detection but rather enhances the robustness and accuracy of anomaly detection. Moreover, UniADC exhibits consistent performance improvement across all settings with the integration of category consistency selection, highlighting its critical role in enhancing the quality of synthesized anomaly samples. More ablation experiments and analysis can be found in the Appendix.
Table 4: Ablation results of UniADC on the MVTec-FS dataset. $\mathcal{L}_{CE}$ denotes the Cross-Entropy loss, and CCSCCS refers to the category consistency selection strategy.
Setting Component I-AUC P-AUC PRO Acc mIoU
Zero-shot $(K_{n}=2)$ w/o $\mathcal{L}_{CE}$ 97.76 96.23 87.10 - -
w/o CCSCCS 95.19 96.02 89.26 66.16 33.15
UniADC 97.09 97.04 92.15 74.74 36.66
Few-shot $(K_{n}=2,K_{a}=1)$ w/o $\mathcal{L}_{CE}$ 97.76 98.33 86.26 - -
w/o CCSCCS 96.84 98.08 88.56 79.91 45.37
UniADC 98.56 98.90 92.48 86.85 51.49
5 Conclusion
This paper integrates image anomaly detection and classification into a unified task and proposes UniADC, an innovative framework capable of anomaly synthesis, detection, and classification. UniADC consists of a training-free controllable inpainting network and a multi-task discriminator. The former can synthesize high-quality anomaly samples of specific categories guided by anomaly priors or anomaly samples, while the latter is trained on these synthesized anomaly samples, achieving unified anomaly detection and classification. Extensive experiments across multiple benchmarks validate the effectiveness and superiority of UniADC. In future work, we aim to further enhance the versatility of our approach by extending its application to diverse domains, such as medical image anomaly detection and classification.
Supplementary Material
Appendix A More Details
This section presents more experimental details, including the experimental setup for the WFDD dataset [3] and a detailed introduction to the anomaly priors provided for UniADC.
A.1 Experimental Setup for WFDD Dataset
The WFDD dataset [3] consists of four types of fabrics: Grey Cloth, Grid Cloth, Pink Flower, and Yellow Cloth. As it was originally designed for the task of anomaly detection, it lacks a detailed categorization of anomaly types. To support our proposed anomaly detection and classification task, we reclassified the anomaly categories, with examples of each anomaly category shown in Fig. S1.
A.2 Anomaly Prior
We employ three types of anomaly priors to characterize various anomaly categories, including shapes, sizes, and textual descriptions. To represent the shapes and sizes of anomalies, we define eight types of anomaly masks: Rectangle, Line, Polygon, Ellipse, Hollow Ellipse, Random Brush, Perlin Noise, and Foreground Mask. The first seven types are used to simulate various local anomalous regions, with each type defined in three sizes: Large, Medium, and Small, as illustrated in Fig. S2. Each mask size corresponds to a specific size range. Except for Perlin noise, all other mask types contain no more than two anomaly regions. Foreground Mask uses the object’s foreground as a mask to synthesize global anomalies. The textual descriptions are used to further define the appearance, attributes, and other relevant characteristics of anomalies. Table S1 presents the list of anomaly priors provided for each anomaly category. For most categories, the anomaly description is a synonym of its name or a combination of its name and the detection object, such as “Foreign matter in the bottle”. For a few anomaly categories, we introduce additional anomaly information, such as the color of the anomalous region, to aid in anomaly classification. We believe that these anomaly priors can be conveniently obtained in real-world scenarios. For example, in the industrial field, the production process and inspection experience of products can offer valuable prior knowledge for anomaly classification.
Figure S1: Examples of different anomaly categories in the WFDD dataset, with anomalous regions marked by red bounding boxes.
Figure S2: Examples of anomaly masks with different types and sizes.
Table S1: Anomaly prior list provided for each anomaly category.
Dataset Image Class Anomaly Category Description Shape Size     Dataset Image Class Anomaly Category Description Shape Size
MVTec-FS Bottle Broken large Badly broken bottle Large     MVTec-FS Screw Manipulated front Damaged screw tip
Broken small Slightly broken bottle Small     Scratch head Damaged screw head
Contamination Foreign matter in the bottle     Scratch neck Damaged screw neck
Cable Bent wire Bending copper wires     Thread side Screw thread stripping
Cable swap Cable swap Foreground     Thread top Damaged screw thread
Cut inner insulation Cracks in the internal insulation     Tile Crack Cracks on the tile
Cut outer insulation Cracks in the external insulation     Glue strip Transparent glue
Missing cable Missing cables Foreground     Gray stroke Gray-black stain Large
Missing wire Missing wires     Oil Oil stain
Poke insulation Puncture-induced holes     Rough Rough white tile Large
Capsule Crack Cracked capsule     Toothbrush Defective Missing bristles, Bending bristle, Foreign matter, A piece of thread
Faulty imprint Font misprint, Font error     Transistor Bent lead Bending metal pins
Poke A tiny hole on the capsule Ellipse Small     Cut lead Broken metal pins
Scratch Slight scratches on the capsule     Damaged case Damaged transistor
Squeeze A capsule deformed by squeezing Large     Misplaced Misplaced transistor Foreground
Carpet Color Red stain, Black stain     Wood Color Red stain, Black stain
Cut A tiny cut in carpet     Hole Hole in wood Ellipse
Hole Hole Ellipse     Liquid Transparent liquid stains
Metal contamination Metal impurities, Metal foreign matter     Scratch Scratches in wood
Thread A piece of thread Line, Hollow Ellipse Small, Medium     Zipper Broken teeth Damage to the zipper
Grid Bent Bending metal grid     Fabric border Border damage of fabric
Broken Broken defect     Fabric interior Interior damage of fabric
Glue Transparent glue     Rough Wear on the zipper
Metal contamination Metal impurities, Metal foreign matter     Split teeth Unzipped zipper
Thread A piece of thread Line, Hollow Ellipse Small, Medium     Squeezed teeth Squeezed zipper
Hazelnut Crack A hazelnut with cracks     WFDD Grey Cloth Contaminated Black stain
Cut A hazelnut with cutting marks     Flecked White spot Ellipse Small
Hole A hazelnut with a hole Ellipse     Line Black linear stain Line Small
Print A hazelnut with printed text     String Fine line Small
Leather Color Red stain, Black stain     Grid Cloth Fold Fabric with fold marks Line
Cut A tiny cut in leather     Flecked White spot Ellipse Small
Fold Fold marks in leather     String Fine line Small
Glue Transparent glue     Pink Flower Hole Hole Ellipse
Poke A puncture-induced hole Ellipse Small     Stain Small stain Small
Metal Nut Bent Cracked metal nut     Tear Zipper-like tear
Color Red stain, Black stain     Yellow cloth Fold Fabric with fold marks Line
Flip Rotating metal nut with motion blur Foreground     Stain Small stain Small
Scratch Scratches on metal nut     String Fine line Small
Pill Color Red stain, Black stain     MTD Blowhole Black hole Ellipse Small
Contamination Foreign matter on the pill     Break Surface crumpling Small
Crack Edge damage of the pill     Crack Surface cracking Small
Faulty imprint Font misprint, Font error     Fray Surface abrasion Large
Pill type Mold growth on the pill Foreground     Uneven Surface unevenness with lighting variation Large
Scratch Scratches on the pill
Appendix B More Results
This section presents additional experimental results, including comparisons with other anomaly synthesis methods, supplementary ablation studies and qualitative results, results under the open-set setting, and per-class metrics.
Table S2: Comparison of UniADC with other anomaly synthesis methods on the MVTec-FS dataset.
@@ Appendix B More Results
B.1 Comparison with Other Anomaly Synthesis Methods
Table S2 presents the performance of UniADC and other anomaly synthesis methods in terms of anomaly detection, classification, and the quality of synthetic anomaly images. For the zero-shot setting, we use CutPaste [19], NSA [29], DRÆM [42], RealNet [46], and AnomalyAny [34] as baseline methods. For the few-shot setting, we use AnoGen [9], DualAnoDiff [15], AnomalyDiffusion [10], as well as the methods without anomaly synthesis (w/o Synthesis) and CutPaste based on masks (CutPasteByMask) as our baseline methods. We use the synthetic anomaly samples from the above methods to train the multi-task discriminator and evaluate their quality using the IS and IC-LPIPS metrics, as in previous works [10, 15]. In both settings, UniADC achieves satisfactory performance and outperforms other methods in multiple aspects. Compared to w/o Synthesis and CutPasteByMask, UniADC shows significant performance improvements, demonstrating the necessity of anomaly sample augmentation and image inpainting.
Table S3: Ablation study of anomaly prior on the MVTec-FS dataset with $K_{n}=2$.
@@ B.1 Comparison with Other Anomaly Synthesis Methods
B.2 Supplementary Ablation Studies
1) Ablation Study on Anomaly Prior. Table S3 investigates the impact of different types of anomaly priors on the performance of UniADC. When anomaly descriptions are unavailable, we use simple combinations of anomaly category names and detection objects as prompts. The experimental results show that anomaly descriptions have the most significant impact on the performance of UniADC, especially in anomaly classification. The introduction of anomaly descriptions enables UniADC to achieve improvements of 4.27% in classification accuracy and 2.78% in mIoU. Compared to anomaly descriptions, shape and size priors are generally difficult to apply to certain types of anomalies. For example, it is hard to predetermine the shapes and sizes of anomaly types such as “Scratch” and “Crack”. Moreover, in our experiments, the size prior yields greater performance improvements for UniADC than the shape prior. When all three types of anomaly priors are applied, UniADC achieves the best overall performance.
Table S4: Ablation study of the noise factor $\gamma$ on the MVTec-FS dataset.

### v2

@@ III-D Implicit-Normal Discriminator
We employ a pre-trained vision encoder to extract multi-scale image features $\{x_{1},x_{2},\cdots,x_{N}\}=\varphi(X)$, and fuse these features into a fine-grained representation through a Feature Fusion Network (FFN) $f=\theta(x_{1},x_{2},\cdots,x_{N})$, where $f\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C}$. As illustrated in Fig. 5, the FFN progressively aggregates intermediate-layer features from high to low semantic levels via a fully convolutional architecture. Concurrently, a text encoder embeds anomaly category names into a semantic space to obtain embeddings $\{g_{1},g_{2},\cdots,g_{Y}\}$, where $g_{y}\in\mathbb{R}^{C}$. The similarity between the anomaly category embedding $g_{y}$ and the feature map $f$ at a spatial location $(h,w)$ is computed as:
Figure 5: Architecture of the Feature Fusion Network.
@@
where $\tau$ represents the anomaly score threshold used to distinguish between normal and anomalous states. We use the anomaly category with the largest pixel area in $S_{c}$ as the image-level anomaly classification result $I_{c}$. If all pixels in $S_{c}$ are classified as normal, then $I_{c}$ is set to 0. We optimize the feature fusion network using Binary Focal loss [45] and Dice loss [46] for anomaly detection, and Cross-Entropy loss for anomaly classification. The overall loss function is:
@@
where $\cdot$$\cdot$] denotes the concatenation operation, and $\lambda$ is a weight hyperparameter. The normal regions in $M_{c}$ are ignored during the computation of $\mathcal{L}_{CE}$. By excluding the normal category, our model is forced to refine the boundaries between various anomaly concepts, thereby enhancing its anomaly recognition performance and classification reliability.
TABLE II: Comparison of UniADC with alternative methods under zero-shot anomaly detection and classification settings. $\mathcal{y}$ denotes methods that use anomaly priors.
$K_{n}\;\;$ Method MVTec-FS [5] MTD [15] WFDD [16]
1 InCTRL [47] 92.36 - - - - 70.34 - - - - 96.66 - - - -
AnomalyDINO [48] 96.12 96.54 89.04 - - 85.32 78.01 74.11 - - 97.42 98.86 87.88 - -
2 InCTRL [47] 93.01 - - - - 72.07 - - - - 97.36 - - - -
AnomalyDINO [48] 96.87 97.14 90.96 - - 86.32 79.27 76.82 - - 97.72 99.13 89.03 - -
4 InCTRL [47] 93.62 - - - - 73.23 - - - - 97.38 - - - -
AnomalyDINO [48] 97.50 97.30 92.21 - - 89.68 83.11 78.61 - - 98.15 99.43 91.54 - -
IV Experiment
To validate the effectiveness of our proposed UniADC framework, we conducted a comprehensive evaluation of its anomaly detection and classification performance on the MVTec-FS [5], MTD [15], and WFDD [16] datasets. This section provides a detailed account of the experimental setup, main experimental results and analysis, as well as a systematic ablation study.
Figure 6: Examples of different anomaly categories in the WFDD dataset, with anomalous regions marked by red bounding boxes.
IV-A Experimental Setup
@@
1) Datasets. We conduct extensive evaluations on three anomaly detection and classification datasets, including MVTec-FS [5], MTD [15], and WFDD [16]. The MVTec-FS dataset is introduced by ZipAF [5] as an extension of the MVTec-AD benchmark [49] for anomaly classification. It contains 15 types of industrial products, each with an average of four anomaly categories. The MTD dataset [15] consists of 1,344 images of magnetic tiles, including five types of anomalies. The subtle differences among these anomalies make the classification task particularly challenging. The WFDD dataset [16] comprises 4,101 images across four fabric types: Grey Cloth, Grid Cloth, Pink Flower, and Yellow Cloth. Since it was originally designed for anomaly detection, it lacked a detailed categorization of anomaly types. To facilitate our proposed anomaly detection and classification task, we reclassified the dataset by grouping identical defects, resulting in an average of three anomaly categories per fabric. Representative examples for each category are illustrated in Fig. 6. Other widely-used anomaly detection datasets, such as VisA [50], are excluded from this work due to the lack of anomaly category labels.
@@
2) Implementation Details. We adopt a consistent experimental setup for both anomaly prior-guided and anomaly sample-guided inpainting. Specifically, we use the DDIM sampler [51] with 1,000 original diffusion steps. The noise factor $\gamma$ is uniformly sampled between 0.4 and 0.6, and the number of accelerated sampling steps is set to 10. For mask generation, we leverage BiRefNet [52] for binary segmentation to ensure that the mask is located within the foreground region. We set the mini-batch size to 32 for category consistency selection, and generate 16 samples with a resolution of $512\times 512$ per anomaly category for discriminator training. For the implicit-normal discriminator, we adopt two experimental settings, namely UniADC(CLIP) and UniADC(DINO), which use the CLIP ViT-L/14 [53] and the DINOv3-based dino.txt [54] as the vision-language backbone of the discriminator, respectively. We set both the anomaly score threshold $\tau$ and the loss weight $\lambda$ to 0.5. The list of anomaly priors is provided in the Supplementary Material.
TABLE III: Comparison of UniADC with alternative methods under few-shot anomaly detection and classification settings.
$K_{n}$ $K_{a}$ Method MVTec-FS [5] MTD [15] WFDD [16]
NAGL+ZipAF [24] 95.60 96.51 92.79 55.92 42.00 84.98 75.45 75.77 36.37 30.89 97.20 98.93 92.40 71.82 45.32
NAGL+ZipAF [24] 96.41 96.56 92.96 57.78 42.34 84.92 77.71 77.72 39.75 31.49 97.47 98.89 92.60 72.40 47.19
NAGL+ZipAF [24] 96.74 97.17 92.80 61.33 45.30 85.77 80.46 78.18 47.21 35.01 97.56 98.90 92.90 76.49 48.85
NAGL+ZipAF [24] 97.14 96.75 93.24 59.49 42.52 85.71 79.49 79.15 42.42 32.07 97.94 99.07 92.72 75.30 49.40
@@
3) Metrics. We use the Area Under the Receiver Operating Characteristic Curve (I-AUC) metric to evaluate image-level anomaly detection performance. For pixel-level anomaly location, we use Pixel-AUROC (P-AUC) and Per Region Overlap (PRO) [55] metrics. In addition, we report the Top-1 classification accuracy (Acc) and Mean Intersection over Union (mIoU) to evaluate image-level and pixel-level anomaly classification performance, respectively. For the “combined” category in MVTec-FS dataset [5], we only evaluate its anomaly detection performance and exclude it from the anomaly classification metrics, as in previous works [5, 6].
@@
4) Baselines. In the zero-shot setting, we adopt InCTRL [47], PatchCore [17], AnomalyDINO [48], PromptAD [8], AnomalyGPT [9], and WinCLIP [7] as baseline methods. For AnomalyGPT [9], we cast classification as a single-choice QA task by prompting the model with candidate anomaly types. For WinCLIP [7], we provide it with anomaly category descriptions and compute the similarity between patch features and anomaly category embeddings, obtaining the anomaly classification results. In the few-shot setting, we combine anomaly detection methods such as PatchCore [17], WinCLIP [7], PromptAD [8], AnomalyCLIP [44], and NAGL [24] with the anomaly classification method ZipAF [5] as our baseline methods. We employ the threshold selection strategy proposed in AnomalyNCD [6] to determine the anomaly score threshold, and fine-tune PromptAD [8] and AnomalyCLIP [44] using few-shot anomaly samples to enhance their anomaly detection performance. To ensure a fair comparison, the visual backbones of CLIP- and DINO-based methods are kept consistent with UniADC(CLIP) and UniADC(DINO), respectively.
Figure 7: Qualitative results of UniADC on the MVTec-FS, MTD, and WFDD datasets under $\text{zero-shot}_{\>(K_{n}=2)}$ and $\text{few-shot}_{\>(K_{n}=2,\,K_{a}=1)}$ settings.
IV-B Experimental Results
@@
1) Anomaly Detection and Classification Results. Tables II and III report the anomaly detection and classification results under zero-shot and few-shot settings, respectively. In the zero-shot setting, most baselines struggle to identify specific anomaly categories due to the absence of real anomaly samples. Although some methods incorporate Vision-Language Models (VLMs) or Large Language Models (LLMs) with anomaly priors, they still fail to achieve satisfactory performance. In contrast, UniADC fully exploits the relevance between anomaly detection and classification, thereby achieving promising performance in both tasks. Compared to other methods, it achieves approximately a 20% improvement in classification accuracy and a 10% improvement in mIoU. Moreover, UniADC(DINO) achieves superior detection and classification accuracy over UniADC(CLIP), highlighting the remarkable performance of DINOv3 [54] in extracting fine-grained image features. In the few-shot setting, UniADC delivers substantial performance gains, even with only one anomaly sample per category, demonstrating its low cost and effective data utilization. Furthermore, it markedly outperforms other two-stage methods, showcasing the benefit of unifying anomaly detection and classification. Qualitative results in Fig. 7 further illustrate its ability to accurately localize and classify anomalous regions under various settings, emphasizing its strong generalization capability and practical applicability.
@@
TABLE IV: Comparison of UniADC with alternative methods on the MVTec-FS dataset using full-shot normal samples, with $K_{a}=1$.
@@
2) Experimental Results with Full-shot Normal Samples. In addition, UniADC can be extended to the setting of full-shot normal samples, with the results provided in Table IV. We use the unsupervised anomaly detection methods PatchCore [17], RD4AD [56], and RealNet [34], as well as the semi-supervised method BGAD [23], in combination with the anomaly classification method ZipAF [5] as our baselines. When sufficient normal samples are available, UniADC can achieve anomaly detection performance comparable to mainstream full-shot methods. Moreover, it significantly improves anomaly classification performance, addressing the limitations of existing approaches in this aspect and demonstrating its unique practical value.
TABLE V: Comparison of UniADC with other anomaly synthesis methods on the MVTec-FS dataset, using DINOv3 as the vision backbone.
@@
Figure 8: Examples of synthetic anomaly samples generated by UniADC under the guidance of anomaly prior and anomaly sample on the MVTec-FS, MTD, and WFDD datasets.
3) Experimental Results on Anomaly Synthesis. Table V presents the performance of UniADC and other anomaly synthesis methods in terms of anomaly detection, classification, and the quality of synthetic anomaly images. For the zero-shot setting, we use CutPaste [30], NSA [31], DRÆM [43], RealNet [34], and AnomalyAny [35] as baseline methods. For the few-shot setting, we use AnoGen [22], DualAnoDiff [39], AnomalyDiffusion [38], as well as the methods without anomaly synthesis (w/o Synthesis) and CutPaste based on masks (CutPasteByMask) as our baseline methods. We use the synthetic anomaly samples from the above methods to train the implicit-normal discriminator and evaluate their quality using the IS and IC-LPIPS metrics, as in previous works [38, 39]. In both settings, UniADC achieves satisfactory performance and outperforms other methods in multiple aspects. Compared to w/o Synthesis and CutPasteByMask, UniADC shows significant performance improvements, demonstrating the necessity of anomaly sample augmentation and image inpainting. Fig. 8 presents synthetic anomaly samples generated by UniADC. By incorporating an inpainting control network and reliable category consistency selection strategies, UniADC can generate high-quality anomaly samples that are mask-consistent, category-aligned, and visually diverse. These properties enable UniADC to generalize well across a wide range of anomaly detection and classification scenarios.
TABLE VI: Comparison of UniADC performance on the Hazelnut class of MVTec-FS dataset under open- and closed-set settings, using DINOv3 as the vision backbone.
Setting I-AUC P-AUC PRO Acc (seen)
Normal Crack Cut Hole AVG
Closed-set 99.63 97.53 91.29 97.50 55.56 62.50 88.89 86.36
Open-set 99.41 99.09 95.21 95.00 44.44 62.50 88.89 83.33
Figure 9: Qualitative results of UniADC under the setting of open-set anomaly detection and classification.
4) Open-set Anomaly Detection and Classification. For open-set anomaly detection, a model is trained on only a subset of anomaly types and is expected to generalize to detect unseen anomalies [57]. We extend this setting to anomaly detection and classification, where a model is designed to detect both seen and unseen anomalies while ensuring correct classification of the seen anomalies. We propose a simple yet effective approach to adapt UniADC for open-set anomaly detection and classification tasks. Specifically, during training, we use unconstrained anomaly synthesis methods (e.g., DRÆM [43]) to generate class-agnostic anomalous images and align them with an additional learnable anomaly embedding. During inference, the learnable embedding is used to match all unseen anomaly types and categorize them as “Other”.
We conducted experiments on the Hazelnut class of the MVTec-FS dataset [5] to validate the effectiveness of our proposed approach. We use “Crack”, “Cut”, and “Hole” as seen anomalies, and “Print” as the unseen anomaly. We adopt the zero-shot setting ($K_{n}=2$), where no prior knowledge about the “Print” category is provided during training. Table VI presents the anomaly detection and classification results of UniADC under both open-set and closed-set settings. In the open-set setting, UniADC achieves anomaly detection and classification performance comparable to the closed-set, demonstrating that it generalizes well to unseen anomalies while avoiding their interference with anomaly classification. Fig. 9 presents qualitative results of UniADC, showing that it ensures accurate detection and classification of seen anomalies, while aligning unseen anomalies with the learnable anomaly embedding to identify them as the “Other” type.
5) More qualitative results and per-class metrics can be found in the Supplementary Material.
TABLE VII: Ablation results of UniADC on the MVTec-FS dataset. $\mathcal{L}_{CE}$ denotes the Cross-Entropy loss, INDIND denotes the Implicit-Normal Discriminator, and CCSCCS refers to the Category Consistency Selection strategy.
Setting Component I-AUC P-AUC PRO Acc mIoU
Zero-shot $(K_{n}=2)$ w/o $\mathcal{L}_{CE}$ 97.76 $(0.67)$ 96.23
$(0.81)$
87.10
$(5.05)$
w/o INDIND 96.72
$(0.37)$
95.70
$(1.34)$
91.44
$(0.71)$
55.91
$(18.83)$
24.99
$(11.67)$
w/o CCSCCS 95.19
$(1.90)$
96.02
$(1.02)$
89.26
$(2.89)$
66.16
$(8.58)$
33.15
$(3.51)$
UniADC 97.09 97.04 92.15 74.74 36.66
Few-shot $(K_{n}=2,K_{a}=1)$ w/o $\mathcal{L}_{CE}$ 97.76
$(0.80)$
98.33
$(0.57)$
86.26
$(6.22)$
w/o INDIND 97.27
$(1.29)$
96.92
$(1.98)$
92.37
$(0.11)$
60.04
$(26.81)$
30.33
$(21.16)$
w/o CCSCCS 96.84
$(1.72)$
98.08
$(0.82)$
88.56
$(3.92)$
79.91
$(6.94)$
45.37
$(6.12)$
UniADC 98.56 98.90 92.48 86.85 51.49
TABLE VIII: Ablation study of anomaly prior on the MVTec-FS dataset with $K_{n}=2$.
@@
IV-C Ablation Study
We conducted extensive experiments to validate the effectiveness of each component in UniADC and its hyperparameter sensitivity. By default, all ablation experiments use DINOv3 [54] as the backbone of the discriminator.
1) Ablation Study on UniADC Components. Table VII investigates the impact of various UniADC components on its performance. When the anomaly classification loss is removed, UniADC degenerates into a standard anomaly detection model. In the zero-shot setting, image-level anomaly detection performance improves slightly, while anomaly localization performance degrades. In the few-shot setting, all metrics decline, indicating that incorporating the anomaly classification task does not compromise anomaly detection but rather enhances the robustness and accuracy of anomaly detection. In the absence of the implicit-normal discriminator, we follow [26] to directly align image patch features with the normal class and $Y$ anomaly classes (totaling $Y+1$ classes), while employing Focal Loss [45] to mitigate pixel distribution imbalance. However, the model experiences a substantial performance decline, with anomaly classification accuracy dropping by approximately 20%. This significant degradation underscores the necessity of implicitly modeling the normal class. Moreover, UniADC exhibits consistent performance improvement across all settings with the integration of category consistency selection, highlighting its critical role in enhancing the quality of synthesized anomaly samples.
2) Ablation Study on Anomaly Prior. Table VIII investigates the impact of different types of anomaly priors on the performance of UniADC. The experimental results show that anomaly descriptions have the most significant impact on the performance of UniADC, especially in anomaly classification. For most anomaly categories, the anomaly description is a synonym of its name or a combination of its name and the detection object, such as “Foreign matter in the bottle”. For a few anomaly categories, we introduce additional anomaly information, such as the color of the anomalous region, to aid in anomaly classification. A comprehensive list of these descriptions is provided in the Supplementary Material. When anomaly descriptions are unavailable, we simply use anomaly category names as prompts. The introduction of anomaly descriptions enables UniADC to achieve improvements of 4.27% in classification accuracy and 2.78% in mIoU. Compared to anomaly descriptions, shape and size priors are generally difficult to apply to certain types of anomalies. For example, it is hard to predetermine the shapes and sizes of anomaly types such as “Scratch” and “Crack”. Moreover, in our experiments, the size prior yields greater performance improvements for UniADC than the shape prior. When all three types of anomaly priors are applied, UniADC achieves the best overall performance.
Figure 10: Examples of synthetic anomaly images generated by UniADC across different noise factors, guided by anomaly prior or anomaly sample, respectively.
TABLE IX: Ablation study of the noise factor on the MVTec-FS dataset.